Skip to end of metadata
Go to start of metadata

11:00 Eastern; 08:00 Project/Pacific Time.



  • Andrés Alejandro Plazas Malagón :
    • Progress on DM-17490 - Getting issue details... STATUS .
      • Posted some plots last week; got feedback from Merlin & Robert.
      • Working on implementing this, then ran into problems with obs_lsst which has delayed this.
        • Potential bug which may require repos rebuilding with new obs_lsst; if this is not sufficient, it might be a bug that needs to get fixed in the v17 release.
    • Further discussions with SAWG, etc
      • Further meeting this week.
      • Planning for upcoming DESC meeting; there will be a SAWG-DM session there.
        • Description from the DESC agenda is: “Brief description of session: Sensor signature removal (ISR) will impact DESC science systematics.  There is a strong need on the DESC side to have  a good understanding of what will be done in DM , and DM has to know where to put its effort on this front. Also the SAWG DESC group developed a good expertise on LSST sensors and has access to test bench & test bench data allowing to probe the sensor signature issue and this information is of interest to DM .  Following an informal series of meeting between  a small set of people from DM and SAWG, we plan to work on specific ISR issues - needs during this session and see how we can further work together.”
  • Unknown User (aguyonnet)
    • Status of paper.
      • Hopefully now has all the MERRA2 data downloaded (yes!).
      • Some bizarre results from MERRA2 data on aerosols about CTIO; mixing latitude and longitude in integration.
        • Now resolved.
      • This week working on refining interpolation above the CTIO site. Seeing oscillations in the aerosol levels, working on understanding why.
        • Estimating altitude for aerosols based on barometric pressure; this may be a source of error.
        • Investigating this on  DM-17834 - Getting issue details... STATUS .
      • Following this work will return to working on the paper.
    • Gave talk on BF effect to EUCLID this morning.
  • Merlin Fisher-Levine
    • Provided a todo list  DM-16699 - Getting issue details... STATUS . Hooray!
    • How far along that list have we progressed? Last week we said another 1.5 weeks of work; is that still plausible?
      • Substantial progress last week.
      • Blocking on John Swinbank getting a new dataset.
    • SAWG meeting last week was “quite upsetting”.
      • “Electronics are a big mess and it's scary”; instability & structure in master biases.
  • John Swinbank
    • Has not followed through on AuxTel schedule with Patrick. Bad John Swinbank — will do that today.
  • Brief discussion